Suppose that the residents of Mexico spend all their income on cauliflower, broccoli, and carrots. In 2017 they buy 110 heads of cauliflower for $220, 80 bunches of broccoli for $100, and 200 carrots for $50. In 2018, they by 50 heads of cauliflower for $200, 50 bunches of broccoli for $110, and 250 carrots for $100. Assume the base year is 2017.

  1. a) What is the CPI in both years? Hints: show steps. (10 marks)
  2. b) What is the inflation rate in 2018? Hints: show steps. (5 marks) 
1
Expert's answer
2021-02-03T02:48:38-0500

a) Let's find the cost of the market basket in the base year (2017):


"Cost_{2017}=\$220+\$100+\$50=\$370."

Then, let's find the cost of the market basket in year 2018:


"Cost_{2018}=\$200+\$110+\$100=\$410."

The Consumer Price Index can be calculated as follows:


"CPI=\dfrac{Cost\ of\ market\ basket\ in\ a\ given\ year}{Cost\ of\ market\ basket\ at\ base\ year}\times100."

Let's calculate CPI in both years:


"CPI_{2017}=\dfrac{Cost_{2017}}{Cost_{2017}}\times100=\dfrac{\$370}{\$370}\times100=100,""CPI_{2018}=\dfrac{Cost_{2018}}{Cost_{2017}}\times100=\dfrac{\$410}{\$370}\times100=111."

b) The % inflation rate in 2018 compared to 2017 can be calculated as follows:


"\%\ Inflation\ rate=\dfrac{CPI_{2018}-CPI_{2017}}{CPI_{2017}}\times100\%,""\%\ Inflation\ rate=\dfrac{111-100}{100}\times100\%=11\%."